# Unlearning Comparator: A Visual Analytics System for Comparative Evaluation of Machine Unlearning Methods

### v1

@@ Abstract
Index Terms:
Machine unlearning, visual analytics, model comparison, model evaluation, data privacy.
@@ I Introduction
We observe that researchers in this emerging field struggle to analyze and understand the behavior of different MU methods. We identify two main hurdles to this challenge. First, despite the recent introduction of MU methods [7, 8, 9, 10, 11], there is still no standardized evaluation protocol for systematically comparing them. Due to the absence of such a protocol, the MU methods have been separately assessed using disparate metrics [12, 13, 14, 15, 16]. This not only hinders direct comparisons among methods but also makes it difficult to understand nuanced trade-offs between the methods. Second, existing evaluations primarily rely on quantitative metrics, which offer limited insight into the behaviors of different methods and fail to uncover the underlying decision-making processes happening in the model, which is insufficient to address the analytic tasks we identified.
For example, one important task in MU research is to evaluate to what extent an MU model satisfies the privacy principle. This principle states that after unlearning, it should be guaranteed that no residual signals of the “forgotten” data remain, thereby thwarting attack [17, 18, 19, 20, 21]. For instance, an attacker might query an online medical diagnostic model to check whether a particular patient’s records still shape its responses, which is an example of membership inference attacks (MIAs). Relying on metrics alone is insufficient for such an evaluation, as such summary metrics would fail to reveal which specific samples remain vulnerable or how model behavior exhibits subtle shifts after unlearning. These limitations call for an interactive approach that allows researchers to simulate attacks and observe how the model responds.
To address the two hurdles, we present Unlearning Comparator, a visual analytics system that enables systematic and multi-foci comparison between MU methods. Our system supports the side-by-side comparison between two models at three aspects: metrics, embeddings, and attacks. By spanning class- and instance-level data as well as output and internal representations, our system aims to show how thoroughly each method removes targeted data without affecting the remainder, highlighting subtle representation changes that aggregate metrics overlook. In our case study and interviews, we demonstrate how this systematic comparison not only clarifies each method’s behavior but also leads to refinements and new designs for a novel MU method. We contribute:
@@ II-A Machine Unlearning Fundamentals
Accuracy. The accuracy principle requires that the forget set should be unlearned without significantly degrading the model’s performance on the retain set. Since deep neural networks have intertwined internal representations, simply dropping certain parts of the representations may harm the overall accuracy. We use four standard accuracy measures: Unlearning Accuracy (UA) measures correctness on the forget set (lower is better) while Retain Accuracy (RA) measures correctness on the retain set (higher is better). If a separate test set is available, the accuracy can also be measured for the unseen data, resulting in Test Unlearning Accuracy (TUA) (lower is better) and Test Retain Accuracy (TRA) (higher is better).
Efficiency. The efficiency principle relates to the computational overhead that an MU method exhibits; for example, building a retrained model can be considered as an MU method but is inefficient as it requires full retraining on the retain set. The efficiency of a method is quantified by its Run Time (RT) represented in seconds.
@@ II-C Distribution-based Approach for Privacy Evaluation
However, evaluating a model’s privacy by simulating such attacks may be less reliable in unlearning scenarios, as these attacks can be easily circumvented; for example, one can deliberately make a model flatten its output scores to increase the entropy and thereby appear resilient to E-MIA. To provide a more rigorous privacy evaluation, we adopt and extend a stronger MIA from prior work [15], which assumes that the attacker compares the unlearned and retrained models’ output distributions. If the unlearned model truly behaves as if it never saw the forget set, its distribution should align with that of the retrained model, providing stronger assurance that no residual data signals remain; in our extension, we incorporate both confidence- and entropy-based attacks under multiple thresholds, yielding a unified Worst-Case Privacy Score (Section IV-A).
II-D Visual Analytics for Machine Learning
@@
Visual analytics has been widely employed to help people develop, interpret, and improve machine learning models [29, 30, 31]. It aims to bridge the gap between complex model behaviors and human understanding, which is essential for building trustworthiness and reliability [32]. For example, Uni-Evaluator [33] supports model evaluation through unified performance analysis across multiple tasks to guide debugging. To help understand internal mechanisms of models, interactive systems that visualize layer-wise activations [34] or architecture‑specific explanations for CNNs [35] and Transformers [36, 37] have also been studied. Furthermore, general-purpose attribution techniques such as Grad-CAM [38] are now widely used to highlight important regions in input images. While these systems and techniques help analyze a single model, they often fall short when there is a need to assess differences between models in a comparative manner, which is also common in machine learning.
Model comparison aims to reveal relative strengths, weaknesses, and behavioral differences between models. Several visual analytics systems have been developed to support comparisons in certain domains of machine learning. Examples include analyzing efficiency-behavior trade-offs in model compression [39], comparing the internal structures of CNN architectures [40], and evaluating behavioral shifts when adapting language models to new tasks [41]. However, these comparison tools are not designed for MU, a nascent field with unique, domain-specific challenges: new concepts such as the distinction of forget and retain sets, the notion of a retrained model as a gold standard, and the need for privacy evaluation.
@@
Overall, although prior VIS4ML systems have effectively supported model understanding across domains, MU introduces distinctive challenges for visual analytics. To address this gap, we present a visual analytics system that enables a unified workflow for the comparative evaluation of MU methods, thereby advancing research in this emerging field.
@@ III Domain Analysis
Our target users are MU researchers who want to compare the performance of different MU methods in terms of the three principles and understand their trade-offs. To identify their needs, we conducted a design study by holding regular discussions over two months with two MU experts to gather insights into their analytical challenges and integrating their feedback throughout the development process. We also surveyed prior literature, including recent benchmarking studies [9, 7, 11], to extract common analysis and evaluation patterns. Synthesizing the insights from both sources allowed us to summarize their goals as follows: (1) evaluate how different MU methods unlearn data by enabling pairwise comparisons as a unified workflow, and (2) leverage these insights to improve MU methods aligned with the three principles.
@@ IV The Unlearning Comparator System
In this section, we elaborate on Unlearning Comparator, a visual analytics system designed for comparative evaluation of MU methods, supporting T1-T5. We chose two datasets, CIFAR-10 [44] and Fashion-MNIST [45], to demonstrate our system, as they are commonly employed in prior MU research. For model architectures, we included two representative architectures: a well-known convolutional architecture ResNet-18 [46] and a more advanced transformer-based architecture, ViT-B/16 [47], pre-trained on ImageNet [48]. While these choices ensure consistency with existing literature, our system is designed to support using other datasets and architectures.
@@ IV-A Designing the Worst-Case Privacy Score
$p_{\text{max}}=\max_{i}p_{i},\quad C(p)=\log(p_{\text{max}})-\log(1-p_{\text{max}})$ (2)
The entropy $H(p)$ is defined with temperature $T=2.0$, selected to better separate model output distributions, as follows:
$p_{i}^{(T)}=\frac{e^{z_{i}/T}}{\sum_{j=1}^{n}e^{z_{j}/T}},\quad H(p)=-\sum_{i=1}^{n}p_{i}^{(T)}\log(p_{i}^{(T)})$ (3)
@@
As shown in Fig. 4, WCPS gradually increases as epochs progress, reflecting that the forget class representations are progressively replaced, and the output distribution converges toward the retrained model’s distribution. However, C-MIA and E-MIA fail to capture this progression. Even at epoch 1, where the distribution has not yet converged (WCPS $\approx$ 0.4), C-MIA prematurely reaches 1.0 (i.e., classifying all forget class samples as non-members). It remains at 1.0 due to its reliance on raw confidence values without log-odds transformation (Eq. 2), which would better separate confidence distributions; this reliance causes incorrect membership inference when confidence drops. E-MIA exhibits even more misleading behavior; it declines as epochs progress because the low entropy from confident misclassification to other classes is misinterpreted as evidence of membership. Notably, the WCPS plateaus around 0.8 rather than reaching 1.0 because FT removes information spanned by the retain classes, struggling to fully erase residual information unique to the forget class [52].
@@ IV-B User Interface
IV-B1 Model Building (the Build stage)
The Model Builder creates unlearned models (Fig. 3A). Users specify the base model to unlearn, the chosen method, and multiple hyperparameter combinations. Models corresponding to all these combinations are automatically generated. To avoid complexity while allowing meaningful adjustments, we limit tunable hyperparameters to epochs, learning rate, and batch size, as these have been shown to play a critical role in prior MU studies [9, 28] and offer sufficient control over the unlearning process. We offer three baseline MU methods from Section II-B (i.e., fine-tuning, random labeling, and gradient ascent).
@@ IV-B1 Model Building (the Build stage)
IV-B2 Model Screening (the Screen stage)
@@
IV-B3 Model Comparison (the Contrast stage)
@@
The Embedding Space view examines the feature space to visually analyze and compare each model’s decision boundary (Fig. 7). We pass each data sample through Model A and Model B up to the penultimate layer to obtain high-dimensional feature vectors, then reduce them to two dimensions and display the resulting scatterplots side-by-side for direct comparison. Because final predictions rely on a linear transformation of these penultimate layer outputs, visualizing them clarifies how the decision boundary is formed. Placing the two scatterplots this way lets users see how an unlearned model’s feature vectors shift from the original model and determine whether they resemble the retrained model’s decision boundary. We use UMAP [62] for dimensionality reduction because it is known to preserve global structure more effectively than other techniques, such as $t$-SNE [63], aiding in locating forget class samples that might scatter unpredictably.
In this view, all training instances are displayed by default (Fig. 2D), but users can activate a highlight mode to focus on specific subsets (Fig. 7A): Target to Forget (all forget class samples), Successfully Forgotten (forget class samples predicted differently from the true label), Not Forgotten (forget class samples still predicted as their true label), or Overly Forgotten (retain class samples misclassified). Users can explore shifts in the feature space to identify areas of interest, such as dense unforgotten clusters (Fig. 7A), and perform instance-level comparisons by linking corresponding instances to track a single sample’s changing predictions and confidences between the two models (Fig. 7B).
IV-B4 Attack Simulation (the Attack stage)
@@
Users can perform various attacks in this view (Fig. 8A). They may select whether to use entropy or confidence and choose different threshold-setting strategies. For example, they can set a threshold to maximize the attack score based on our defined FPR and FNR, select a threshold with higher overall success, or apply a common threshold for both Model A and Model B to achieve the highest attack score under identical conditions. Ultimately, across two attack metrics, threshold directions, and strategies, the worst-case scenario—where the model is most vulnerable—defines the WCPS, and the user can check exactly which configuration leads to that worst case via a dedicated button.
@@ IV-C Implementation Details
Our implementation consists of a React.js [^1] https://react.dev frontend that provides visualizations powered by D3.js [^2] https://d3js.org and follows the Shadcn/UI [^3] https://ui.shadcn.com design system, and a FastAPI [^4] https://fastapi.tiangolo.com backend that manages model evaluation, unlearning operations, and data handling. The system provides pre-computed results for system-provided models stored as JSON files, with their checkpoints automatically downloaded from Hugging Face repositories [^5] https://huggingface.co/jaeunglee/resnet18-cifar10-unlearning, https://huggingface.co/Yurim0507/resnet18-fashionmnist-unlearning, https://huggingface.co/Yurim0507/vit-base-16-cifar10-unlearning upon backend initialization. When users configure unlearning parameters and initiate execution through the UI, the frontend sends API calls to trigger real-time unlearning on the backend. Upon completion, the backend saves both model checkpoints and metrics for visualization as JSON files. The JSON is then transmitted to the frontend for rendering. Additionally, users can add custom MU methods by implementing Python hooks with minimal boilerplate code; once registered, these methods appear in the UI alongside existing MU methods. The source code is publicly available at https://github.com/gnueaj/Machine-Unlearning-Comparator.
@@ V Case Study
To evaluate if and how our system assists users in understanding the behavior of MU methods, we conducted a case study in collaboration with two experts introduced in Section III. To the best of our knowledge, this study presents the first comparative visual analysis of prominent MU methods to enhance their transparency and interpretability [65]. The case study consists of two stages. In the first analysis stage, the experts freely used our system to compare the behavior of five MU methods: three baseline methods (FT, RL, and GA) and two state-of-the-art MU methods (SCRUB and SalUn) as described in Section II-B. To examine whether the insights from the analysis stage could lead to actual improvements in MU performance, we conducted a second, improvement stage, in which the experts developed a novel MU method based on their findings from the analysis stage. For the case study, we employed a ResNet-18 architecture on the CIFAR-10 dataset, designating various classes (e.g., “airplane”, “automobile,” or “deer”) as the forget class one at a time.

### v3

@@ I Introduction
We observe that researchers in this emerging field struggle to analyze and understand the behavior of different MU methods. We identify two main hurdles to this challenge. First, despite the recent introduction of MU methods [36, 21, 6, 11, 23], there is still no standardized evaluation protocol for systematically comparing them. Due to the absence of such a protocol, the MU methods have been separately assessed using disparate metrics [27, 2, 15, 60, 41]. This not only hinders direct comparisons among methods but also makes it difficult to understand nuanced trade-offs between the methods. Second, existing evaluations primarily rely on quantitative metrics, providing limited insight into the behaviors of different methods and failing to uncover the underlying decision-making processes happening in the model, which is insufficient to address the analytic tasks we identified.
One critical principle in MU evaluation is the privacy, which requires that no residual signals of the forgotten data remain, thereby thwarting attacks [55, 22, 12, 8, 33]. For instance, an attacker might query an online medical diagnostic model to check whether a particular patient’s records still shape its responses, which is an example of membership inference attacks (MIAs). Relying on metrics alone is insufficient for such an evaluation, as such summary metrics would fail to reveal which specific samples remain vulnerable or how model behavior exhibits subtle shifts after unlearning. Consequently, there is a need for an interactive diagnostic approach that enables researchers to verify granular privacy risks through attack simulations, instead of being limited to aggregate metrics.
To address the two hurdles, we present Unlearning Comparator, a visual analytics system that enables systematic and multi-foci comparison between MU methods. Our system supports the side-by-side comparison between two models at three aspects: metrics, embeddings, and attacks. By spanning class- and instance-level data as well as output and internal representations, our system aims to show how thoroughly each method removes targeted data without affecting the remainder, highlighting subtle representation changes that aggregate metrics overlook. In our case study and interviews, we demonstrate how this systematic comparison enables a human-in-the-loop process, where researchers combine algorithmic insights with visual evidence to integrate ideas from existing methods into a novel MU method design. We contribute:
@@ II-A Machine Unlearning Fundamentals
Accuracy. The accuracy principle requires that the forget set should be unlearned without significantly degrading the model’s performance on the retain set. Since deep neural networks have intertwined internal representations, simply dropping certain parts of the representations may harm the overall accuracy. We use four standard accuracy measures: Unlearning Accuracy (UA) measures correctness on the forget set (lower is better), while Retain Accuracy (RA) measures correctness on the retain set (higher is better). If a separate test set is available, the accuracy can also be measured for the unseen data, resulting in Test Unlearning Accuracy (TUA) (lower is better) and Test Retain Accuracy (TRA) (higher is better).
Efficiency. The efficiency principle relates to the computational overhead that an MU method exhibits; for example, building a retrained model can be considered as an MU method but is inefficient as it requires full retraining on the retain set. The efficiency of a method is quantified by its Run Time (RT), measured in seconds.
@@ II-C Distribution-based Approach for Privacy Evaluation
However, evaluating a model’s privacy by simulating such attacks may be less reliable in unlearning scenarios, as these attacks can be easily circumvented; for example, one can deliberately make a model flatten its output scores to increase the entropy and thereby appear resilient to E-MIA. To provide a more rigorous privacy evaluation, we adopt and extend a stronger MIA from prior work [60], which assumes that the attacker compares the unlearned and retrained models’ output distributions. If the unlearned model truly behaves as if it never saw the forget set, its distribution should align with that of the retrained model, providing stronger assurance that no residual data signals remain; in our extension, we incorporate both confidence- and entropy-based attacks under multiple thresholds, yielding a unified Worst-Case Privacy Score (see Section IV-A).
II-D Visualization for Model Understanding and Comparison
@@
Visual analytics has been widely employed to help people develop, interpret, compare, and improve machine learning models [58, 30, 62]. To bridge the gap between complex model behaviors and human understanding [9], prior work has introduced systems ranging from unified performance evaluation [10] to internal mechanism inspection via layer-wise activations [37] or architecture-specific explanations for CNNs [63] and Transformers [69, 13]. Furthermore, attribution techniques such as Grad-CAM [51] are now widely used to highlight important regions in input images. While these systems and techniques help analyze a single model, they often fall short when assessing differences between models in a comparative manner, which is also common in machine learning.
Existing visual analytics systems for model comparison provide workflows for revealing relative strengths, weaknesses, and behavioral differences between models. For example, Manifold [72] supports pairwise comparison to localize performance gaps, and ModelWise [45] enables multi-model comparison to surface when and why models disagree for iterative improvement. Other domain-specific systems compare models for compression trade-offs [5], architectural differences of CNNs [68], or behavioral shifts in language model adaptation [53]. However, these generic workflows are inadequate for the unique design space of MU. While comparison typically aims to identify a “better” model under fixed ground-truth labels, MU requires behavioral verification: assessing whether an unlearned model aligns with a “gold-standard” retrained model that varies with the forget set. Moreover, the unlearning process involves opposing optimization objectives: removing information in the forget set while preserving it in the retain set. Systems treating datasets holistically cannot effectively guide the trade-off analysis required for this dichotomy. Finally, privacy in MU is a behavioral property assessed through adversarial verification (e.g., MIAs), an aspect outside standard model comparison workflows.
@@
To address this distinct design space, we present a visual analytics system that formalizes these elements into a unified workflow for the comparative evaluation of MU methods, thereby advancing research in this emerging field.
@@ III Domain Analysis
Our target users are MU researchers who want to compare the performance of different MU methods in terms of the three principles and understand their trade-offs. To identify their needs, we conducted a design study by holding weekly discussions over two months with two MU experts (one professor specializing in ML security including the privacy concerns of MU and one graduate student with years of experience in MU research) to gather insights into their analytical challenges and integrating their feedback throughout the development process. We also surveyed prior literature, including recent benchmarking studies [6, 36, 23], to extract common analysis and evaluation patterns. Synthesizing the insights from both sources allowed us to summarize their goals as follows: (1) evaluate how different MU methods unlearn data by enabling pairwise comparisons as a unified workflow, and (2) leverage these insights to improve MU methods aligned with the three principles.
@@ IV The Unlearning Comparator System
In this section, we elaborate on Unlearning Comparator, a visual analytics system designed for comparative evaluation of MU methods, supporting T1-T5. We chose two datasets, CIFAR-10 [39] and Fashion-MNIST [66], to demonstrate our system, as they are commonly employed in prior MU research. For model architectures, we included two representative architectures: a well-known convolutional architecture ResNet-18 [29] and a more advanced transformer-based architecture, ViT-B/16 [20], pre-trained on ImageNet [16]. While these choices ensure consistency with existing literature, our current system targets ten-class image classification datasets and corresponding model architectures. We discuss extensions beyond this setting in Section VII.
@@ IV-A Designing the Worst-Case Privacy Score
$p_{\text{max}}=\max_{i}p_{i},\quad C(p)=\log(p_{\text{max}})-\log(1-p_{\text{max}}).$ (2)
The entropy $H(p)$ is defined with temperature $T=2.0$, chosen empirically to better spread the entropy values for visual inspection, as follows:
$p_{i}^{(T)}=\frac{e^{z_{i}/T}}{\sum_{j=1}^{n}e^{z_{j}/T}},\quad H(p)=-\sum_{i=1}^{n}p_{i}^{(T)}\log(p_{i}^{(T)}).$ (3)
@@
As shown in Fig. 4, WCPS gradually increases as epochs progress, reflecting that the forget class representations are progressively replaced, and the output distribution converges toward the retrained model’s distribution. However, C-MIA and E-MIA fail to capture this progression. Even at epoch 1, where the distribution has not yet converged (WCPS $\approx$ 0.4), C-MIA prematurely reaches 1.0 (i.e., classifying all forget class samples as non-members). It remains at 1.0 due to its reliance on raw confidence values without log-odds transformation (Eq. 2), which would better separate confidence distributions; this reliance causes incorrect membership inference when confidence drops. E-MIA exhibits even more misleading behavior; it declines as epochs progress because the low entropy from confident misclassification to other classes is misinterpreted as evidence of membership. Notably, the WCPS plateaus around 0.8 rather than reaching 1.0 because FT removes information spanned by the retain classes, struggling to fully erase residual information unique to the forget class [18]; this motivates stronger or complementary strategies beyond FT to achieve more complete erasure.
@@ IV-B User Interface
IV-B1 Model Building (the Build stage, T1)
The Model Builder creates unlearned models (Fig. 3A). Users specify the base model to unlearn, the chosen method, and multiple hyperparameter combinations. Models corresponding to all these combinations (epochs $\times$ learning rates $\times$ batch sizes) are automatically generated. To avoid complexity while allowing meaningful adjustments, we limit tunable hyperparameters to epochs, learning rate, and batch size, as these have been shown to play a critical role in prior MU studies [6, 40] and offer sufficient control over the unlearning process. We offer three baseline MU methods from Section II-B (i.e., fine-tuning, random labeling, and gradient ascent).
@@ IV-B1 Model Building (the Build stage, T1)
IV-B2 Model Screening (the Screen stage, T1)
@@
IV-B3 Model Comparison (the Contrast stage, T2T4)
@@
The Embedding Space view visualizes the feature space to analyze and compare each model’s decision boundary (Fig. 7). We pass each data sample through Model A and Model B up to the penultimate layer to obtain high-dimensional feature vectors, then reduce them to two dimensions and display the resulting scatterplots side-by-side. Because final predictions are a linear transformation of these penultimate representations, this view reveals how an unlearned model’s features shift from the original model and whether they align with the retrained model. For example, after successful unlearning, the forget class embeddings should align with the retrained model, whereas persistent forget class clusters or distorted retain class clusters (e.g., density changes) can indicate under- or over-forgetting. We use UMAP [44] for dimensionality reduction because it is known to preserve global structure more effectively than other techniques, such as $t$-SNE [61], aiding in locating forget class samples that might scatter unpredictably.
In this view, all training instances are displayed by default (Fig. 2D), but users can activate a highlight mode to focus on specific subsets (Fig. 7A): Target to Forget (all forget class samples), Successfully Forgotten (forget class samples predicted differently from the true label), Not Forgotten (forget class samples still predicted as their true label), or Overly Forgotten (retain class samples misclassified). Users can explore shifts in the feature space to identify areas of interest, such as dense unforgotten clusters (Fig. 7A-right), and perform instance-level comparisons by linking corresponding instances to track a single sample’s changing predictions and confidences between the two models (Fig. 7B).
IV-B4 Attack Simulation (the Attack stage, T5)
@@
Users can perform various attacks in this view (Fig. 8A). They may select whether to use entropy or confidence and choose different threshold-setting strategies. For example, they can set a threshold to maximize the attack score based on our defined FPR and FNR, select a threshold with higher overall success, or apply a common threshold for both Model A and Model B yielding the highest attack score to compare their vulnerability under the same threshold setting. Ultimately, across two attack metrics, threshold directions, and strategies, the worst-case scenario—where the model is most vulnerable—defines the WCPS, and the user can automatically apply the configuration via a dedicated button.
@@ IV-C Implementation Details
Our implementation consists of a React.js frontend that provides visualizations powered by D3.js and follows the Shadcn/UI design system, and a FastAPI backend that manages model evaluation, unlearning operations, and data handling. The system provides pre-computed results for system-provided models stored as JSON files, with their checkpoints automatically downloaded from Hugging Face repositories [^1] https://huggingface.co/jaeunglee/resnet18-cifar10-unlearning, https://huggingface.co/Yurim0507/resnet18-fashionmnist-unlearning, https://huggingface.co/Yurim0507/vit-base-16-cifar10-unlearning upon backend initialization. When users configure unlearning parameters and initiate execution through the UI, the frontend sends API calls to trigger real-time unlearning on the backend. On a single NVIDIA A6000 GPU, unlearning typically takes from several seconds to about 5 minutes depending on the method and architecture. The subsequent evaluation pipeline (e.g., CKA computation, dimensionality reduction) requires approximately 30 seconds. Upon completion, the backend saves both model checkpoints and metrics for visualization as JSON files. The JSON is then transmitted to the frontend for rendering. Additionally, users can add custom MU methods by implementing Python hooks with minimal boilerplate code; once registered, these methods appear in the UI alongside existing MU methods. The source code is publicly available at https://github.com/gnueaj/Machine-Unlearning-Comparator.
@@ V Case Study
To evaluate if and how our system assists users in understanding the behavior of MU methods, we conducted a case study in collaboration with two experts introduced in Section III. To the best of our knowledge, this study presents the first comparative visual analysis of prominent MU methods to enhance their transparency and interpretability [54]. The case study consists of two stages. In the first analysis stage, the experts freely used our system to compare the behavior of five MU methods: three baseline methods (FT, RL, and GA) and two state-of-the-art MU methods (SCRUB and SalUn) as described in Section II-B. To examine whether the insights from the analysis stage could lead to actual improvements in MU performance, we conducted a second, improvement stage, in which the experts developed a novel MU method based on their findings from the analysis stage. For the case study, we employed a ResNet-18 architecture on the CIFAR-10 dataset, designating various classes (e.g., “airplane,” “automobile,” or “deer”) as the forget class one at a time.